### v2

# From Limited Labels to Open Domains:An Efficient Learning Method for Drone-view Geo-Localization

Abstract
Traditional supervised drone-view geo-localization (DVGL) methods heavily depend on paired training data and encounter difficulties in learning cross-view correlations from unpaired data. Moreover, when deployed in a new domain, these methods require obtaining the new paired data and subsequent retraining for model adaptation, which significantly increases computational overhead. Existing unsupervised methods have enabled to generate pseudo-labels based on cross-view similarity to infer the pairing relationships. However, geographical similarity and spatial continuity often cause visually analogous features at different geographical locations. The feature confusion compromises the reliability of pseudo-label generation, where incorrect pseudo-labels drive negative optimization. Given these challenges inherent in both supervised and unsupervised DVGL methods, we propose a novel cross-domain invariant knowledge transfer network (CDIKTNet) with limited supervision, whose architecture consists of a cross-domain invariance sub-network (CDIS) and a cross-domain transfer sub-network (CDTS). This architecture facilitates a closed-loop framework for invariance feature learning and knowledge transfer. The CDIS is designed to learn cross-view structural and spatial invariance from a small amount of paired data that serves as prior knowledge. It endows the shared feature space of unpaired data with similar implicit cross-view correlations at initialization, which alleviates feature confusion. Based on this, the CDTS employs dual-path contrastive learning to further optimize each subspace while preserving consistency in a shared feature space. Extensive experiments demonstrate that CDIKTNet achieves state-of-the-art performance under full supervision compared with those supervised methods, and further surpasses existing unsupervised methods in both few-shot and cross-domain initialization.
Index Terms:
drone-view geo-localization, structural invariance learning, spatial invariance learning, knowledge transfer.
I Introduction
Cross-view geo-localization has emerged as a promising visual localization technique [shore2024bev]. Early efforts primarily focus on retrieving ground-view images to geo-referenced satellite imagery for ground-view geo-localization [Cai_2019_ICCV]. With the widespread adoption of drones in complex scenarios such as emergency response [Liu_Multi] and urban surveys [Li_PVT++], the demand for autonomous localization in GPS-denied environments has been steadily increasing. This has driven more research interests in drone-view geo-localization (DVGL) [deuser2023sample4geo, wang2024multiple]. However, the dynamic viewpoints of drones result in scale distortions and non-rigid spatial transformations compared with satellite imagery. This pronounced structural and spatial heterogeneity presents challenges for DVGL. In recent years, benefiting from previous research efforts [wu2024camp, chen2024multi], supervised DVGL methods with paired data have made notable progress. These paired data often require manual labeling, which provides an accurate pairing relationship between satellite imagery and drone images of the same location.
[…]
Figure 2: CDIKTNet vs. State-of-the-Art Unsupervised Methods. (a) Spatial continuity causes visually similar patterns from different locations. Unsupervised methods often encounter feature confusion, which results in inaccurate pseudo-labels and drives negative optimization. (b) CDIKTNet leverages a small amount of paired data at initialization to learn cross-view invariance as prior knowledge. Based on this, it obtains a reliable shared feature space and further employs dual-path contrastive learning to optimize each subspace.
However, as illustrated in Fig. 1 (A(a)) and (B(a)), existing supervised methods [shen2023mccg, du2024ccr, xia2024enhancing] strictly rely on a large amount of paired data to learn cross-view correlations. The absence of sufficient paired data leads to a dramatic decline in model performance. When the supervised ratio drops to 0, or in other words, when there is no paired data and all the data are unpaired, the existing supervised methods become ineffective. We mark this situation with a red X in the figures. Furthermore, dynamic drone mission environments often lead to domain shifts, which in turn require repeated efforts to re-establish cross-view correlations before deployment in new domains. It hinders the applicability in scenarios where pairing relationships are difficult to obtain and cross-domain adaptation is required. As shown in Fig. 1 (B(a)), the new domain where all data are unpaired cannot directly employ the trained model from the source domain. Therefore, researchers have increasingly focused on unsupervised methods for DVGL. Recent work [11010141] constructed a sparse 3D point cloud from multi-view drone images and rendered pseudo-satellite images with near-orthographic views via 3D Gaussian splatting to realize geometric transformations of the drone images. These synthetic images are then matched against real satellite imagery to enable unsupervised training. However, this method is sensitive to 3D reconstruction and pose estimation errors, and it is difficult to bridge the inherent cross-view domain gaps. In contrast, another work [li2024learning] utilizes a frozen pre-trained backbone and a learnable adapter to align features from different views in shared feature space. The shared feature space serves as a unified latent representation in which drone and satellite views can be directly aligned. Cross-view pseudo-labels are generated automatically based on feature similarity as supervised signals to guide training. This method avoids explicit geometric transformations and aims to intrinsically learn cross-view correlations. Nevertheless, as shown in Fig. 2 (a), when visually similar patterns appear at different geographical locations, the resulting feature confusion in the shared feature space often leads to incorrect pseudo-label generation. This ultimately drives the model toward negative optimization.
In this paper, we propose an efficient learning network with limited supervision for DVGL that mitigates the adverse effects of inaccurate pseudo-labels caused by spatial ambiguity, called the cross-domain invariant knowledge transfer network (CDIKTNet). Specifically, the CDIKTNet consists of two core components, including the cross-domain invariance sub-network (CDIS) and the cross-domain transfer sub-network (CDTS). The CDIS utilizes a small amount of paired data to learn latent cross-view structural and spatial invariance, which serves as prior knowledge. As illustrated in Fig. 2(b), it endows the shared feature space of unpaired data with similar implicit cross-view correlations at initialization, which alleviates feature confusion. Based on this, CDTS introduces dual-path contrastive learning to independently optimize subspaces under different views. Despite being optimized separately, these subspaces maintain alignment in the shared feature space due to the shared invariance prior. The main contributions can be summarized as follows.
  • We propose a cross-domain invariant knowledge transfer network, which can leverage few-shot guided cross-view invariance learning to further explore latent cross-view correlations within unpaired data. It reduces reliance on strictly paired data and enables direct transfer to new domains without requiring any prior pairing information.
  • The CDIS is designed to learn the structural and spatial invariance by integrating channel-wise local and global structural information, and performing joint interaction learning along the different dimensions of the features. It is utilized to construct a shared feature space with better consistency and compactness for subsequent unsupervised learning.
  • The CDTS is built on the shared cross-view invariance prior, which enables unpaired data to form initial cross-view correlations in the shared feature space. On this basis, dual-path contrastive learning is employed to perform clustering and assign pseudo-labels within subspaces, which are then used as supervised signals to optimize the representations of each view.
  • Three experimental settings are designed to encompass a spectrum of learning paradigms that span fully supervised learning and limited-supervision learning. In the context of limited supervised learning, the proportion of paired data and unpaired data is allowed for random variation. Specifically, the cross-domain initialization setting provides an effective learning paradigm for new domains without requiring any prior pairing information. Extensive experiments demonstrate that CDIKTNet outperforms existing supervised methods under full supervision and surpasses existing unsupervised methods in both few-shot initialization and cross-domain initialization setting.
Notably, as illustrated in Fig. 1(A(b)) and (B(b)), compared with supervised methods, CDIKTNet remains effective at varying levels of supervision and can be transferred to new domains without requiring paired data. Furthermore, as shown in Fig. 2, compared with unsupervised methods, CDIKTNet performs initial feature space optimization using limited paired data. Meanwhile, it adopts dual-path contrastive learning to optimize each subspace, which reduces feature confusion.
The remainder of this paper is organized as follows. Section II systematically reviews previous research. In Section III, the proposed method is presented in detail. The experimental results are reported and analyzed in Section LABEL:result. Finally, conclusions are outlined in Section LABEL:conclusion.
II Related Work
II-A Supervised Cross-view Geo-localization
Cross-view geo-localization (CVGL) is widely recognized as an image retrieval problem based on heterogeneous visual data. However, challenges arise from geometric-semantic heterogeneity and feature distribution shifts across different viewpoints. In traditional CVGL sub-tasks, such as ground-view geo-localization [Liu_2019_CVPR, Hu, Shi_Yu_Liu_Zhang_Li_2020], many researchers have employed explicit geometric alignment methods [shi2019spatial, zhang2023cross, 10601183] to enforce structural consistency at varying viewpoints, while others [regmi2019bridging, Toker_2021_CVPR] have used GANs [goodfellow2014generative] to map source domain images to the target domain. Moreover, in drone-view geo-localization, which is another core sub-task in CVGL, PCL [tian2021uav] attempted to align aerial and satellite images using perspective projection transformation. It also employed GAN to synthesize satellite-style images in order to reduce spatial discrepancies. But it was difficult to handle non-rigid deformations caused by dynamic UAV viewpoints and remained constrained by the synthetic-to-real domain gap. Recently, researchers [ge2024multibranch, lv2024direction] have begun to model latent geometric structures or spatial consistency to enhance the robustness of DVGL. Nevertheless, these methods often focused solely on optimizing geometric structures or spatial representations while neglecting their synergistic relationships. This results in limited generalization under scale variations and spatial shifts in the DVGL. Furthermore, these methods still rely on large amounts of paired data for training, which restricts their applicability in open environments.
Inspired by the aforementioned supervised learning methods, our method utilizes a small amount of paired data for initialization , and further exploits structural and spatial consistency under limited supervision to learn robust and generalizable cross-view representations. Compared with these supervised methods, our method significantly reduces the reliance on paired data while achieving superior performance.
II-B Unsupervised Cross-Domain Retrieval
In recent years, cross-domain image retrieval based on the supervised methods has made significant progress [10314802, 10812860, 9963608, 10855601]. However, the reliance on large-scale annotated data has limited its applicability, prompting researchers to explore unsupervised cross-domain retrieval methods [10109197, 9783116, 9891821]. Current methods [Chen_2021_ICCV, Xuan_2021_CVPR, dai2022cluster] utilized powerful feature backbone networks for feature initialization, employing clustering algorithms [ester1996density] to generate pseudo-labels and feature queues that provide stable negative samples for optimizing feature learning. However, these methods are primarily designed for homogeneous data with strong correlations, and their effectiveness tends to degrade significantly when applied to heterogeneous scenarios. To address the limitations, some researchers [yang2022augmented, yang2024shallow] have proposed cross-modal person retrieval methods based on dual-path contrastive learning. Nevertheless, these methods still depend on pre-trained person retrieval models for feature extraction during initialization.
To better learn cross-view correlations, we adopt a pseudo-label-based training scheme using unpaired data, where pseudo-labels are generated through clustering, similar to the methods discussed previously. In contrast, we initialize the model with a small amount of paired data. It leads to a more compact and discriminative feature space and enables more reliable pseudo-label generation. Notably, although unsupervised DVGL can also be considered an unsupervised cross-domain retrieval task, existing methods remain limited and have been discussed in the introduction. We refrain from repeating the details here.
III Proposed Method
Problem Formulation: Given a drone-view image set $\mathcal{X}^{d}=\{x_{i}^{d}\}_{i=1}^{N}$ and a satellite-view image set $\mathcal{X}^{s}=\{x_{j}^{s}\}_{j=1}^{M}$, where each drone image $x_{i}^{d}$ is paired with a corresponding satellite image $x_{j}^{s}$ as a positive pair $(x_{i}^{d},x_{j}^{s})$, the objective is to learn a shared feature space $\mathcal{V}$ that minimizes the distance between positive pairs while maximizing the distance between negative pairs. Existing methods heavily rely on a large amount of paired data as positive samples for training to achieve high performance. In contrast, our method fundamentally differs by mitigating this dependency. The following sections provide a detailed explanation of our method.
Figure 3: Pipeline Overview. The proposed network pipeline consists of two sub-networks. The cross-domain invariance sub-network is designed to learn feature representations that exhibit structural invariance and spatial invariance. The cross-domain transfer sub-network utilizes the invariant knowledge learned by the cross-domain invariance sub-network as a feature representation space to further explore the latent correlations of cross-view from unpaired data.
III-A Cross-Domain Invariance Sub-network
The CDIS lies in the complementarity between structural and spatial invariance. In DVGL, variations in viewpoint and scale are significant. Therefore, it is necessary to learn the structural and spatial features that are invariant to varying viewpoints and scales. This structural invariance consists of local geometric information, like corners, edges, and contours, and global structural information, such as spatial layout and domain-invariant topologies. Additionally, spatial invariance is derived from the scale-aware contextual cues.
III-A1 Structural Invariant Learning (SIFL) Block
Given a pair of cross-view images $\{x_{i}^{d},x_{j}^{s}\}$, high-dimensional feature representations are first extracted using the ConvNeXt-Base [liu2022convnet] backbone network. The extracted features are represented as $\mathbf{f}_{v}^{u}\in\mathbb{R}^{C\times H\times W}$, where $u\in\{d,s\}$. Specifically, $v$ is defined as follows: when $u=d$, $v=i$ and when $u=s$, $v=j$. To effectively leverage structural features at different viewpoints, the dimensions $H\times W$ are flattened into a single dimension $L$ to obtain the features $\mathbf{f}_{v}^{ul}\in\mathbb{R}^{C\times L}$. $\mathbf{f}_{v}^{ul}$ serves as the input to the SIFL block.
Subsequently, a depthwise separable $3\times 3$ convolution $\varphi^{3}_{q}$ is used to mine local structural priors, while residual connections preserve the original topological information. After normalization $\mathcal{N}(\cdot)$, the feature representation $\mathbf{f}_{v+}^{u}$ can be obtained. To further enhance the capability of structural invariance feature learning, we proposed a channel-wise refinement (CWR) module. Specifically, we project $\mathbf{f}_{v+}^{u}$ into the latent space through two cascaded linear transformation layers $\mathcal{I}(\cdot)$ to learn high-order interactions between channels. A subsequent $1\times 1$ convolution $\psi^{1}_{q}$ is applied to explicitly capture local structural features.
Following this, a gating mechanism $\mathcal{G}(\cdot)$ consists of a depthwise separable $3\times 3$ convolution $\varphi_{k}^{3}$ branch and a linear transformation branch $L(\cdot)$, which is performed the channel-wise selective enhancement via a sigmoid activation. Then, the residual connection and normalization $\mathcal{N}(\cdot)$ are employed to fuse cross-layer information to enhance representation. This process can be formulated as
$\mathbf{f}_{v+}^{u}=\mathcal{N}(\varphi^{3}_{q}(\mathbf{f}_{v}^{ul})+\mathbf{f}_{v}^{ul})$ (1)
$\mathbf{f}_{v+}^{u+}=\mathcal{N}(\mathcal{G}(\psi^{1}_{q}(\mathcal{I}(\mathbf{f}_{v+}^{u}))))+\varphi^{3}_{q}(\mathbf{f}_{v}^{ul})+\mathbf{f}_{v}^{ul}.$ (2)
To further enhance local geometric information, we adopt the same feature processing procedure as in Eq. 1 to extract the feature $\mathbf{f}_{v++}^{u+}$. Although $\mathbf{f}_{v++}^{u+}$ contains rich local geometric information, it still lacks global structural features. Therefore, we replace the CWR module described previously with an MLP $\mathcal{P}(\cdot)$ combined with the squeeze-and-excitation (SE) attention [hu2018squeeze]. In this process, the global average pooling $\mathcal{A}(\cdot)$ is first applied within the SE attention to extract the global information at the channel level. The channel information is then compressed through two linear transformation layers $\mathcal{I}^{+}(\cdot)$, followed by the channel attention weights calculated using the sigmoid function $\mathcal{F}$, which are then combined with the global channel information obtained through the MLP. This further enhances the expression of global channel-level information. Finally, a residual connection is used to fuse the rich local channel-level information from the original features with the enhanced global information. This process can be formulated as
$\mathbf{f}_{v+}^{u++}=\mathcal{N}(\varphi^{3}_{q}(\mathbf{f}_{v+}^{u+})+\mathbf{f}_{v+}^{u+})$ (3)
$\mathbf{f}_{v++}^{u++}\!\!=\!\!\mathcal{F}(\mathcal{I}^{+}(\mathcal{A}(\mathbf{f}_{v++}^{u+})))\times\mathcal{P}(\mathbf{f}_{v+}^{u++})+\varphi^{3}_{q}(\mathbf{f}_{v+}^{u+})+\mathbf{f}_{v+}^{u+}.$ (4)
III-A2 Spatial Invariant Enhancement Learning (SIEL) Module
Spatial information is particularly crucial for DVGL. To enhance spatial invariance, inspired by [shen2023mccg], we design a dual-path spatial invariance enhancement learning module to mine spatial information from different dimensions.
In the spatial invariance enhancement learning module, we preserve the dimensions, namely height $H$ and weight $W$ of the input features. By permuting the input features $\mathbf{f}_{v}^{u}\in\mathbb{R}^{C\times H\times W}$ along the channel-height (C-H) and channel-width (C-W) dimensions, we obtain $\mathbf{f}_{v}^{hu}\in\mathbb{R}^{H\times C\times W}$ and $\mathbf{f}_{v}^{wu}\in\mathbb{R}^{W\times H\times C}$, respectively. This allows the module to capture multi-perspective dependencies from different dimensions. Subsequently, global feature representations $\mathbf{f}_{v2}^{hu}\in\mathbb{R}^{2\times H\times W}$ and $\mathbf{f}_{v2}^{wu}\in\mathbb{R}^{2\times H\times W}$ are generated by concatenating the results of max pooling and average pooling applied to the permuted features. A $3\times 3$ depthwise separable convolution $\varphi^{3}_{v}$ is then employed to obtain a more compact feature representation. Subsequently, a strong cross-interaction across the C-H and C-W is performed, followed by separate $1\times 1$ convolution $\psi^{1}_{k}$ to further enhance the spatial invariance of the feature representation. Finally, the fused features are normalized using a sigmoid function $\mathcal{F}$ to generate attention weights, which are multiplied with the original feature maps and restored to the original dimensions $\mathbb{R}^{C\times H\times W}$, yielding the spatially enhanced feature $\mathbf{f}_{v+}^{hu}\in\mathbb{R}^{C\times H\times W}$ and $\mathbf{f}_{v+}^{wu}\in\mathbb{R}^{C\times H\times W}$ as
$\mathbf{f}_{v+}^{hu}=\mathcal{F}(\psi^{1}_{k}(\varphi^{3}_{v}(\mathbf{f}_{v2}^{hu}))+\psi^{1}_{k}(\mathbf{f}_{v2}^{wu}))\times\mathbf{f}_{v}^{u}$ (5)
$\mathbf{f}_{v+}^{wu}=\mathcal{F}(\psi^{1}_{k}(\varphi^{3}_{v}(\mathbf{f}_{v2}^{wu}))+\psi^{1}_{k}(\mathbf{f}_{v2}^{hu}))\times\mathbf{f}_{v}^{u}.$ (6)
III-A3 Multi-level Optimization
[…]
$\mathcal{L}_{1}=\lambda_{1}\mathcal{L}_{MSE}+\lambda_{2}\mathcal{L}_{CE}+\lambda_{3}\mathcal{L}_{infoNCE}$ (7)
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are the coefficients to control the relative importance of the loss terms.
Remark 1
The CDIS lies in dual optimization of structural and spatial invariance. On one hand, it integrates local and global structural information along the channel dimension and is guided by $\mathcal{L}_{MSE}$ to effectively constrain the feature search space, which enhances the structural representations. On the other hand, CDIS retains the original spatial dimensions and performs joint interaction learning along C-H and C-W, where $\mathcal{L}_{CE}$ is employed to further mine scale-aware contextual cues. This structure–spatial collaborative modeling enables the CDIS to achieve rapid convergence and learn highly discriminative representations, even with limited paired data.
III-B Cross-Domain Transfer Sub-network
The CDTS further explores cross-view correlations based on the enhanced consistency and compactness established by the CDIS. Specifically, the CDTS incorporates dual-path contrastive learning tailored explicitly for DVGL.
III-B1 Cross-View Feature Memory Initialization
As shown in Fig. 3, at the beginning of each training epoch, we initialize cross-view feature instances using the invariant knowledge learned by the CDIS and perform clustering to generate pseudo-labels. This process aims to mitigate the issue of low-quality pseudo-label initialization caused by suboptimal feature extraction. Subsequently, we construct cluster memories for each view, specifically for drone and satellite perspectives denoted as $\phi_{k}^{d}$ and $\phi_{l}^{s}$, by computing the average feature representations of drone and satellite instances within each cluster. This process can be formulated as
$\phi_{k}^{d}=\frac{1}{|\mathcal{H}_{k}^{d}|}\sum_{\mathbf{f}_{i}^{d}\in\mathcal{H}_{k}^{d}}\mathbf{f}_{i}^{d},\phi_{l}^{s}=\frac{1}{|\mathcal{H}_{l}^{s}|}\sum_{\mathbf{f}_{j}^{s}\in\mathcal{H}_{l}^{s}}\mathbf{f}_{j}^{s}$ (8)
where $\mathbf{f}_{i}^{d}$ and $\mathbf{f}_{j}^{s}$ are drone and satellite instance features, respectively. $i=1,\cdots,N$, $j=1,\cdots,M$. $\mathcal{H}_{k}^{d}$ and $\mathcal{H}_{l}^{s}$ denote the $k$-th cluster set in drone and $l$-th cluster satellite view, respectively. $|\cdot|$ indicates the number of instances per cluster.
III-B2 Cross-View Feature Memory Updating
During each training iteration, the cluster memories using a momentum updating strategy are represented as
$\quad\phi_{k}^{d(\omega)}\leftarrow\alpha\phi_{k}^{d(\omega-1)}+(1-\alpha)q_{d}$ (9)
$\quad\phi_{k}^{s(\omega)}\leftarrow\alpha\phi_{k}^{s(\omega-1)}+(1-\alpha)q_{s}$ (10)
where $\alpha$ is momentum updating factor and $\omega$ is iteration number. $q_{d}$ and $q_{s}$ represent the given query features of the drone and the satellite, respectively.
III-B3 Cross-View Contrastive Loss
Given the drone and satellite query $q_{d}$ and $q_{s}$, we compute the contrastive loss for the drone view and the satellite view as
$\mathcal{L}^{d}=-\log\frac{\exp(q_{d}\cdot\phi^{d}_{+}/\tau)}{\sum_{k=0}^{K}\exp(q_{d}\cdot\phi^{d}_{k}/\tau)}$ (11)
$\mathcal{L}^{s}=-\log\frac{\exp(q_{s}\cdot\phi^{s}_{+}/\tau)}{\sum_{l=0}^{L}\exp(q_{s}\cdot\phi^{s}_{l}/\tau)}$ (12)
where $\tau$ denotes the temperature hyper-parameter. Drone cluster memories $\phi^{d}_{k}$ and satellite cluster memories $\phi^{s}_{l}$ are utilized as feature vectors at the cluster level to compute the distances between the query instance $q_{d(s)}$ and all clusters. Here, $\phi^{d}_{+}$ and $\phi^{s}_{+}$ correspond to the positive cluster memory associated with $q_{d}$ and $q_{s}$, respectively.