### v4

# From Limited Labels to Open Domains:An Efficient Learning Method for Drone-view Geo-Localization

Abstract
Traditional supervised drone-view geo-localization (DVGL) methods heavily depend on paired training data and encounter difficulties in learning cross-view correlations from unpaired data. Moreover, when deployed in a new domain, these methods require obtaining the new paired data and subsequent retraining for model adaptation, which significantly increases computational overhead. Existing unsupervised methods have enabled to generate pseudo-labels based on cross-view similarity to infer the pairing relationships. However, geographical similarity and spatial continuity often cause visually analogous features at different geographical locations. The feature confusion compromises the reliability of pseudo-label generation, where incorrect pseudo-labels often drive negative optimization. Given these challenges inherent in both supervised and unsupervised DVGL methods, we propose a novel cross-domain invariant knowledge transfer network (CDIKTNet) with limited supervision, whose architecture consists of a cross-domain invariance sub-network (CDIS) and a cross-domain transfer sub-network (CDTS). This architecture facilitates a closed-loop framework for invariance feature learning and knowledge transfer. The CDIS is designed to learn cross-view structural and spatial invariance from a small amount of paired data that serves as prior knowledge. It endows the shared feature space of unpaired data with implicit cross-view correlations at initialization, which alleviates feature confusion. Based on this, CDTS employs dual-path contrastive learning to further optimize each subspace while preserving consistency in a shared feature space. Extensive experiments demonstrate that CDIKTNet achieves state-of-the-art performance under full supervision compared with those supervised methods, and further surpasses existing unsupervised methods in both few-shot initialization and cross-domain initialization settings.
I Introduction
[…]
Cross-view geo-localization has emerged as a promising visual localization technique [39]. Early efforts primarily focus on retrieving ground-view images to geo-referenced satellite-view images for ground-view geo-localization [3]. With the widespread adoption of drones in complex scenarios such as emergency response [29] and urban surveys [22], the demand for autonomous localization in GPS-denied environments has been steadily increasing. This has driven more research interests in drone-view geo-localization (DVGL) [9, 45]. However, the dynamic viewpoints of drones result in scale distortions and non-rigid spatial transformations compared with satellite-view images. This pronounced structural and spatial heterogeneity presents challenges for DVGL. In recent years, benefiting from previous research efforts [47, 5], supervised DVGL methods with paired data have made notable progress. These paired data often require manual labeling, which provides an accurate pairing relationship between satellite-view and drone-view images of the same location.
Figure 2: CDIKTNet vs. State-of-the-Art Unsupervised Methods. (a) Spatial continuity causes visually similar patterns from different locations. Unsupervised methods often encounter feature confusion, which results in inaccurate pseudo-labels and drives negative optimization. (b) CDIKTNet leverages a small amount of paired data at initialization to learn cross-view invariance as prior knowledge. Based on this, it obtains a reliable shared feature space and further employs dual-path contrastive learning to optimize each subspace.
However, as illustrated in Fig. 1 (A(a)) and (B(a)), existing supervised methods [36, 10, 48] strictly rely on a large amount of paired data to learn cross-view correlations. The lack of sufficient paired data leads to a dramatic decline in model performance. When the supervised ratio drops to 0, or in other words, when there is no paired data and all the data are unpaired, the existing supervised methods become ineffective. We mark this situation with a red X in the figures. Furthermore, dynamic drone mission environments often lead to domain shifts, which in turn require repeated efforts to re-establish cross-view correlations before deployment in new domains. It hinders the applicability in scenarios where pairing relationships are difficult to obtain and cross-domain adaptation is required. As shown in Fig. 1 (B(a)), the new domain where all data are unpaired cannot directly employ the trained model from the source domain. Therefore, researchers have increasingly focused on unsupervised methods for DVGL. Recent work [23] constructed a sparse 3D point cloud from multi-view aerial images and rendered pseudo-satellite images with near-orthographic views via 3D Gaussian splatting to realize geometric transformations of drone-view images. These synthetic images are then matched against real satellite-view images to enable unsupervised training. However, this method is sensitive to 3D reconstruction and pose estimation errors, and it is difficult to bridge the inherent cross-view domain gaps. In contrast, another work [24] utilizes a frozen pre-trained backbone and a learnable adapter to align features from different viewpoints in a shared feature space. The shared feature space serves as a unified latent representation in which drone and satellite views can be directly aligned. Cross-view pseudo-labels are generated automatically based on feature similarity as supervised signals to guide training. This method avoids explicit geometric transformations and aims to intrinsically learn cross-view correlations. Nevertheless, as shown in Fig. 2 (a), when visually similar patterns appear at different geographical locations, the resulting feature confusion in the shared feature space often leads to incorrect pseudo-label generation. This ultimately drives the model toward negative optimization.
In this paper, we propose an efficient learning method with limited supervision for DVGL that mitigates the adverse effects of inaccurate pseudo-labels caused by feature confusion, called the cross-domain invariant knowledge transfer network (CDIKTNet). Specifically, the CDIKTNet consists of two core components, including the cross-domain invariance sub-network (CDIS) and the cross-domain transfer sub-network (CDTS). The CDIS utilizes a small amount of paired data to learn latent cross-view structural and spatial invariance, which serves as prior knowledge. As illustrated in Fig. 2(b), it endows the shared feature space of unpaired data with similar implicit cross-view correlations at initialization to alleviate feature confusion. Based on this, the CDTS introduces dual-path contrastive learning to independently optimize subspaces under different views. Despite being optimized separately, these subspaces maintain alignment in the shared feature space due to the shared invariance prior. The main contributions can be summarized as follows.
  • We propose a cross-domain invariant knowledge transfer network, which can leverage few-shot guided cross-view invariance learning to further explore latent cross-view correlations within unpaired data. It reduces reliance on strictly paired data and enables direct transfer to new domains without requiring any prior pairing information.
  • The CDIS is designed to learn the structural and spatial invariance by integrating channel-wise local and global structural information, and performing joint interaction learning along the different dimensions of the features. It is utilized to construct a shared feature space with better compactness for subsequent unsupervised learning.
  • The CDTS is built on the shared cross-view invariance prior, which enables unpaired data to form initial cross-view correlations in the shared feature space. On this basis, dual-path contrastive learning is employed to perform clustering and assign pseudo-labels within subspaces, which are then used as supervised signals to optimize the representations of each view.
  • Three experimental settings are designed to encompass a spectrum of learning paradigms that span fully supervised learning and limited-supervision learning. In the context of limited supervised learning, the proportion of paired data and unpaired data is allowed for random variation. Specifically, the cross-domain initialization setting provides an effective learning paradigm for new domains without requiring any prior pairing information. Extensive experiments demonstrate that CDIKTNet outperforms existing supervised methods under full supervision and surpasses existing unsupervised methods in both few-shot initialization and cross-domain initialization setting.
Notably, as illustrated in Fig. 1(A(b)) and (B(b)), compared with supervised methods, CDIKTNet remains effective at varying levels of supervision and can be transferred to new domains without requiring paired data. Furthermore, as shown in Fig. 2, compared with unsupervised methods, CDIKTNet performs initial feature space optimization using limited paired data. Meanwhile, it adopts dual-path contrastive learning to optimize each subspace, which reduces feature confusion.
The remainder of this paper is organized as follows. Section II systematically reviews previous research. In Section III, the proposed method is presented in detail. The experimental results are reported and analyzed in Section IV. Finally, conclusions are outlined in Section V.
II Related Work
II-A Supervised Cross-view Geo-localization
Cross-view geo-localization (CVGL) is widely recognized as an image retrieval problem based on heterogeneous visual data. However, challenges arise from geometric-semantic heterogeneity and feature distribution shifts across different viewpoints. In traditional CVGL sub-tasks, such as ground-view geo-localization [27, 18, 38], many researchers have employed explicit geometric alignment methods [37, 55, 1] to enforce geometric consistency at varying viewpoints, while others [35, 43] have used GANs [16] to map source domain images to the target domain, which aims to reduce appearance discrepancies and maintain style consistency. Moreover, in drone-view geo-localization, which is another sub-task in CVGL, PCL [42] attempted to align drone-view and satellite-view images using perspective projection transformation. It also employed GAN to synthesize satellite-style images in order to reduce cross-view discrepancies and enhance both geometric and style consistency. But it was difficult to handle non-rigid deformations caused by dynamic UAV viewpoints and remained constrained by the synthetic-to-real domain gap. Moreover, such explicit feature consistency mappings fail to capture intrinsic cross-view invariance that differs from these environmentally sensitive features [5, 6]. Recently, researchers [15, 32] have begun to model latent structures or spatial invariance to enhance the robustness of DVGL. Nevertheless, these methods often focused solely on optimizing structures or spatial representations while neglecting their synergistic relationships. This results in limited generalization under scale variations and spatial shifts in the DVGL. Furthermore, these methods still rely on large amounts of paired data for training, which restricts their applicability in open environments.
Inspired by the aforementioned supervised learning methods, our method utilizes a small amount of paired data for initialization, and further exploits structural and spatial invariance under limited supervision to learn robust and generalizable cross-view representations. Compared with these supervised methods, our method significantly reduces the reliance on paired data while achieving superior performance.
II-B Unsupervised Cross-view Retrieval
Supervised methods for cross-view retrieval have achieved remarkable progress [13, 20]. Early cross-view retrieval tasks, exemplified by person re-identification [28, 54, 2, 53], along with the current ground-view geo-localization [18, 38] and drone-view geo-localization tasks [5, 15] that have been attracting extensive attention, effectively establish the instance-level feature consistency through aligning directly based on feature similarity in supervised scenarios. However, the reliance on large-scale annotated data has limited its applicability. Fortunately, recent advances in domain adaptation [58, 30, 25] and semi-unsupervised learning [40, 19, 11] have demonstrated that robust representations can be learned even under weak or missing labels. This research trend provides important insights for extending cross-view retrieval toward unsupervised settings and has motivated researchers to explore unsupervised cross-view retrieval methods [46, 41, 34]. Current methods [4, 49, 8] utilized powerful feature backbone networks for feature initialization, employing clustering algorithms [12] to generate pseudo-labels and feature queues that provide stable negative samples for optimizing feature learning. However, these methods are primarily designed for homogeneous data with strong correlations, and their effectiveness tends to degrade significantly when applied to heterogeneous scenarios. To address the limitations, some researchers [52, 50] have proposed cross-modal person retrieval methods based on dual-path contrastive learning. Nevertheless, these methods still depend on pre-trained person retrieval models for feature extraction during initialization.
To better learn cross-view correlations, we adopt a pseudo-label-based training scheme using unpaired data, where pseudo-labels are generated through clustering in accordance with the methods described above. In contrast, the model is initialized with a small amount of paired data. It leads to a more compact and discriminative feature space and enables more reliable pseudo-label generation. Notably, although unsupervised DVGL can also be considered an unsupervised cross-domain retrieval task, existing methods remain limited and have already been discussed in the Introduction. We refrain from repeating the details here.
III Proposed Method
Problem Formulation. Given a drone-view image set $\mathcal{X}^{d}=\{x_{i}^{d}\}_{i=1}^{N}$ and a satellite-view image set $\mathcal{X}^{s}=\{x_{j}^{s}\}_{j=1}^{M}$, where each drone-view image $x_{i}^{d}$ is paired with a corresponding satellite-view image $x_{j}^{s}$ as a positive pair $(x_{i}^{d},x_{j}^{s})$, the objective is to learn a shared feature space $\mathcal{V}$ that minimizes the distance between positive pairs while maximizing the distance between negative pairs. Existing methods heavily rely on a large amount of paired data as positive samples for training to achieve high performance. In contrast, our method fundamentally differs by mitigating this dependency. The following sections provide a detailed explanation of our method.
Figure 3: Pipeline Overview. The proposed network pipeline consists of two sub-networks. The cross-domain invariance sub-network is designed to learn feature representations that exhibit structural invariance and spatial invariance. The cross-domain transfer sub-network utilizes the invariant knowledge learned by the cross-domain invariance sub-network as a feature representation space to further explore the latent cross-view correlations from unpaired data.
III-A Cross-Domain Invariance Sub-network
The CDIS lies in the complementarity between structural and spatial invariance. In DVGL, variations in viewpoint and scale are significant. Therefore, it is necessary to learn the structural and spatial features that are invariant to varying viewpoints and scales. This structural invariance consists of local geometric information, like corners, edges, and contours, and global structural information, such as spatial layout and domain-invariant topologies. Additionally, spatial invariance is derived from the scale-aware contextual cues.
III-A1 Structural Invariant Learning (SIFL) Block
Given a pair of cross-view images $\{x_{i}^{d},x_{j}^{s}\}$, high-dimensional feature representations are first extracted using the ConvNeXt-Base [31] backbone network. The extracted features are represented as $\mathbf{f}_{v}^{u}\in\mathbb{R}^{C\times H\times W}$, where $u\in\{d,s\}$. Specifically, $v$ is defined as follows: when $u$ is $d$ , $v$ is ii and when $u$ is $s$, $v$ is $j$. To effectively leverage structural features at different viewpoints, the dimensions $H\times W$ are flattened into a single dimension $L$ to obtain the features $\mathbf{f}_{v}^{ul}\in\mathbb{R}^{C\times L}$. $\mathbf{f}_{v}^{ul}$ serves as the input to the SIFL block.
Subsequently, a depthwise separable $3\times 3$ convolution $\varphi^{3}_{q}$ is used to mine local structural priors, while residual connections preserve the original topological information. After normalization $\mathcal{N}(\cdot)$, the feature representation $\mathbf{f}_{v+}^{u}$ can be obtained. To further enhance the capability of structural invariance feature learning, we proposed a channel-wise refinement (CWR) module. Specifically, we project $\mathbf{f}_{v+}^{u}$ into the latent space through two cascaded linear transformation layers $\mathcal{I}(\cdot)$ to learn high-order interactions between channels. A subsequent $1\times 1$ convolution $\psi^{1}_{q}$ is applied to explicitly learn local structural features.
Following this, a gating mechanism $\mathcal{G}(\cdot)$ consists of a depthwise separable $3\times 3$ convolution $\varphi_{k}^{3}$ branch and a linear transformation branch $L(\cdot)$, which is performed the channel-wise selective enhancement via a sigmoid activation. Then, the residual connection and normalization $\mathcal{N}(\cdot)$ are employed to fuse cross-layer information to enhance representation. This process can be formulated as
$\mathbf{f}_{v+}^{u}=\mathcal{N}\left(\varphi^{3}_{q}(\mathbf{f}_{v}^{ul})+\mathbf{f}_{v}^{ul}\right)$ (1)
$\mathbf{f}_{v+}^{u+}=\mathcal{N}\left(\mathcal{G}\left(\psi^{1}_{q}\left(\mathcal{I}\left(\mathbf{f}_{v+}^{u}\right)\right)\right)\right)+\varphi^{3}_{q}\left(\mathbf{f}_{v}^{ul}\right)+\mathbf{f}_{v}^{ul}$ (2)
To further enhance local geometric information, we adopt the same feature processing procedure as in Eq. 1 to extract the feature $\mathbf{f}_{v++}^{u+}$. Although $\mathbf{f}_{v++}^{u+}$ contains rich local geometric information, it still lacks global structural features. Therefore, we replace the CWR module described previously with an MLP $\mathcal{P}(\cdot)$ combined with the squeeze-and-excitation (SE) attention [17]. In this process, the global average pooling $\mathcal{A}(\cdot)$ is first applied within the SE attention to extract the global information at the channel level. The channel information is then compressed through two linear transformation layers $\mathcal{I}^{+}(\cdot)$, followed by the channel attention weights calculated using the sigmoid function $\mathcal{F}$, which are then combined with the global channel information obtained through the MLP. This further enhances the expression of global channel-level information. Finally, a residual connection is used to fuse the rich local channel-level information from the original features with the enhanced global information. This process can be formulated as
$\mathbf{f}_{v++}^{u+}=\mathcal{N}\left(\varphi^{3}_{q}\left(\mathbf{f}_{v+}^{u+}\right)+\mathbf{f}_{v+}^{u+}\right)$ (3)
$\mathbf{f}_{v++}^{u++}\!\!=\!\!\mathcal{F}\left(\mathcal{I}^{+}\left(\mathcal{A}\left(\mathbf{f}_{v++}^{u+}\right)\right)\right)\times\mathcal{P}\left(\mathbf{f}_{v++}^{u+}\right)+\varphi^{3}_{q}\left(\mathbf{f}_{v+}^{u+}\right)+\mathbf{f}_{v+}^{u+}$ (4)
III-A2 Spatial Invariant Enhancement Learning (SIEL) Module
Spatial information is particularly crucial for DVGL. To enhance spatial invariance, inspired by [36], we design a dual-path spatial invariance enhancement learning module to mine spatial information from different dimensions.
In the spatial invariance enhancement learning module, we preserve the dimensions, namely height $H$ and weight $W$ of the input features. By permuting the input features $\mathbf{f}_{v}^{u}\in\mathbb{R}^{C\times H\times W}$ along the channel-height (C-H) and channel-width (C-W) dimensions, we obtain $\mathbf{f}_{v}^{hu}\in\mathbb{R}^{H\times C\times W}$ and $\mathbf{f}_{v}^{wu}\in\mathbb{R}^{W\times H\times C}$, respectively. This allows the module to capture multi-perspective dependencies from different dimensions. Subsequently, global feature representations $\mathbf{f}_{v2}^{hu}\in\mathbb{R}^{2\times C\times H}$ and $\mathbf{f}_{v2}^{wu}\in\mathbb{R}^{2\times W\times C}$ are generated by concatenating the results of max pooling and average pooling applied to the permuted features. A $3\times 3$ depthwise separable convolution $\varphi^{3}_{v}$ and a separate $1\times 1$ convolution $\psi^{1}_{k}$ are then employed to obtain a more compact feature representation. Subsequently, cross-interaction across the C-H and C-W is performed to further enhance the spatial invariance of the feature representation. Finally, the fused features are normalized using a sigmoid function $\mathcal{F}$ to generate attention weights, which are multiplied with the original feature maps and restored to the original dimensions $\mathbb{R}^{C\times H\times W}$, yielding the spatially enhanced feature $\mathbf{f}_{v+}^{hu}\in\mathbb{R}^{C\times H\times W}$ and $\mathbf{f}_{v+}^{wu}\in\mathbb{R}^{C\times H\times W}$ as
$\mathbf{f}_{v+}^{hu}$ $=\mathcal{F}\left(\psi^{1}_{k}\left(\varphi^{3}_{v}\left(\mathbf{f}_{v2}^{hu}\right)\right)+\psi^{1}_{k}\left(\varphi^{3}_{v}\left(\mathbf{f}_{v2}^{wu}\right)\right)\right)\times\mathbf{f}_{v}^{hu}$ (5)
$\mathbf{f}_{v+}^{wu}$ $=\mathcal{F}\left(\psi^{1}_{k}\left(\varphi^{3}_{v}\left(\mathbf{f}_{v2}^{wu}\right)\right)+\psi^{1}_{k}\left(\varphi^{3}_{v}\left(\mathbf{f}_{v2}^{hu}\right)\right)\right)\times\mathbf{f}_{v}^{wu}$ (6)
III-A3 Multi-level Optimization
In the CDIS, the features $\mathbf{f}_{v++}^{u++}$ obtained from the SIFL, which denotes as $\mathbf{f}_{i++}^{d++}$ for drone-view intermediate features and $\mathbf{f}_{j++}^{s++}$ for satellite-view intermediate features, are optimized using the mean squared error loss $\mathcal{L}_{MSE}$ [48]. Furthermore, to enhance structural representation in the spatial invariance enhancement learning module, the features $\mathbf{f}_{i}^{ad}$ and $\mathbf{f}_{j}^{as}$ are obtained by applying global average pooling to $\mathbf{f}_{i}^{d}$ and $\mathbf{f}_{j}^{s}$ , respectively. The feature $\mathbf{f}_{i}^{ad}$ combined with $\mathbf{f}_{i+}^{hd}$ and $\mathbf{f}_{i+}^{wd}$ is passed through a classifier module [36] to obtain $\mathbf{f}_{i+}^{cd}$, as well as $\mathbf{f}_{j+}^{cs}$ obtained in the same manner. Subsequently, $\mathbf{f}_{i+}^{cd}$ and $\mathbf{f}_{j+}^{cs}$ are optimized with cross-entropy loss $\mathcal{L}_{CE}$ to enhance feature discrimination. In addition, to further enhance the feature representations, $\mathbf{f}_{i}^{ad}$ and $\mathbf{f}_{j}^{as}$ undergo independent optimization using InfoNCE loss $\mathcal{L}_{InfoNCE}$ [33]. The final objective is to minimize the sum of all losses. It can be formulated as
$\mathcal{L}_{1}=\lambda_{1}\mathcal{L}_{MSE}+\lambda_{2}\mathcal{L}_{CE}+\lambda_{3}\mathcal{L}_{infoNCE}$ (7)
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ serve as loss coefficients to balance the relative contributions of individual loss terms.
Remark 1
The CDIS lies in dual optimization of structural and spatial invariance. On one hand, it integrates local and global structural information along the channel dimension and is guided by $\mathcal{L}_{MSE}$ to effectively constrain the feature search space, which enhances the structural representations. On the other hand, the CDIS retains the original spatial dimensions and performs joint interaction learning along C-H and C-W, where $\mathcal{L}_{CE}$ is employed to further mine scale-aware contextual cues. This structure–spatial collaborative modeling enables CDIS to achieve rapid convergence and learn highly discriminative representations, even with limited paired data.
III-B Cross-Domain Transfer Sub-network
The CDTS further explores cross-view correlations based on the enhanced compactness established by CDIS. Specifically, the CDTS incorporates dual-path contrastive learning tailored explicitly for DVGL.
III-B1 Cross-View Feature Memory Initialization
As shown in Fig. 3, at the beginning of each training epoch, cross-view feature instances are initialized using the invariant knowledge learned by CDIS and then clustering is performed to generate pseudo-labels. It can mitigate the issue of low-quality pseudo-label initialization caused by suboptimal feature extraction. Subsequently, cluster memories are constructed for the drone-view and satellite-view. These memories $\{\phi_{k}^{d}\}_{k=1}^{K}$ and $\{\phi_{l}^{s}\}_{l=1}^{L}$ are obtained by computing the average feature representation of all drone or satellite instances within each cluster, where $K$ and $L$ denote the total number of drone-view clusters and satellite-view clusters, respectively. $\phi_{k}^{d}$ and $\phi_{l}^{s}$ serve as the cluster centroids for the drone-view and satellite-view features, respectively. This process can be formulated as
$\phi_{k}^{d}=\frac{1}{|\mathcal{H}_{k}^{d}|}\sum_{\mathbf{f}_{i}^{d}\in\mathcal{H}_{k}^{d}}\mathbf{f}_{i}^{d},\phi_{l}^{s}=\frac{1}{|\mathcal{H}_{l}^{s}|}\sum_{\mathbf{f}_{j}^{s}\in\mathcal{H}_{l}^{s}}\mathbf{f}_{j}^{s}$ (8)
where $\mathbf{f}_{i}^{d}$ and $\mathbf{f}_{j}^{s}$ are drone and satellite instance features, respectively. $i=1,\cdots,N$, $j=1,\cdots,M$. $\mathcal{H}_{k}^{d}$ and $\mathcal{H}_{l}^{s}$ denote the $k$-th cluster in drone-view and $l$-th cluster in satellite-view, respectively. $|\cdot|$ indicates the number of instances per cluster.
Remark 2
In selecting the clustering algorithm, we refer to several unsupervised cross-view retrieval works [[8, 51] and adopt DBSCAN [12] for clustering. Unlike partition-based methods such as K-means [26], DBSCAN utilizes a density-based method, which handles clusters of arbitrary shapes and noisy data. It does not require a predefined number of clusters, nor does it force all points to be assigned to a cluster. Additionally, DBSCAN can identify and exclude noise points, which ensures the selection of valid samples even in cross-view challenges.
III-B2 Cross-View Feature Memory Updating
During each training iteration, the cluster memories using a momentum updating strategy are represented as
$\quad\phi_{k}^{d(\omega)}\leftarrow\alpha\phi_{k}^{d(\omega-1)}+(1-\alpha)q_{d}$ (9)
$\quad\phi_{l}^{s(\omega)}\leftarrow\alpha\phi_{l}^{s(\omega-1)}+(1-\alpha)q_{s}$ (10)
where $\alpha$ is the momentum factor and $\omega$ is the iteration number. $q_{d}$ and $q_{s}$ represent the given query features of the drone and the satellite, respectively.
III-B3 Cross-View Contrastive Loss
Given the drone and satellite query $q_{d}$ and $q_{s}$, we compute the contrastive loss for the drone view and the satellite view as
$\mathcal{L}^{d}=-\log\frac{\exp(q_{d}\cdot\phi^{d}_{+}/\tau)}{\sum_{k=1}^{K}\exp(q_{d}\cdot\phi^{d}_{k}/\tau)}$ (11)
$\mathcal{L}^{s}=-\log\frac{\exp(q_{s}\cdot\phi^{s}_{+}/\tau)}{\sum_{l=1}^{L}\exp(q_{s}\cdot\phi^{s}_{l}/\tau)}$ (12)
where $\tau$ denotes the temperature hyper-parameter. $\phi^{d}_{k}$ (or $\phi^{s}_{l}$ ) is utilized as a feature vector at the cluster level to compute the distances between the query instance $q_{d}$ (or $q_{s}$) and all clusters. Here, $\phi^{d}_{+}$ and $\phi^{s}_{+}$ correspond to the positive cluster memory associated with $q_{d}$ and $q_{s}$ in $\{\phi_{k}^{d}\}_{k=1}^{K}$ and $\{\phi_{l}^{s}\}_{l=1}^{L}$, respectively. In practice, for a given query feature, e.g., $q_{d}$ or $q_{s}$, the cluster memory that is assigned in the same cluster, namely $\phi^{d}_{+}$ or $\phi^{s}_{+}$, is designated as the positive sample.
Remark 3
Benefiting from CDIS, instances $\mathbf{f}_{i}^{d}$ and $\mathbf{f}_{j}^{s}$ exhibit stronger compactness within their respective subspaces and cross-view correlation in shared feature space. This also makes $\phi_{k}^{d}$ and $\phi_{l}^{s}$ more representative and an initial alignment in the shared feature space. On this basis, $\mathcal{L}^{d}$ and $\mathcal{L}^{s}$ further guide each instance feature towards its corresponding $\phi_{k}^{d}$ and $\phi_{l}^{s}$ within the subspace, thereby indirectly facilitating cross-view alignment in the shared feature space.
The CDTS is optimized jointly with $\mathcal{L}^{d}$ and $\mathcal{L}^{s}$. Thus, the final optimization for CDIKTNet is denoted as $\mathcal{L}\!=\!\mathcal{L}_{1}\!+\!\mathcal{L}^{d}\!+\!\mathcal{L}^{s}$.
IV Experimental Results
IV-A Datasets and Evaluation Protocol
CDIKTNet is evaluated on two challenging benchmarks for DVGL. University-1652 dataset [57] consists of drone, satellite, and ground-view images from 1,652 locations in 72 universities. Its training set includes 701 buildings from 33 universities, while the test set comprises 951 buildings from 39 universities. SUES-200 dataset [59] comprises 200 locations, which are split into 120 for training and 80 for testing. Each location provides a satellite-view image and drone-view images taken at four different altitudes, which encompass diverse environments such as parks, lakes, and buildings. We employ Recall@K (R@K) and Average Precision (AP) as evaluation metrics. Additionally, Ground Truth Ratio (GT ratio) denotes the proportion of paired data within the entire experimental dataset.
IV-B Implementation Details
During training and testing, all input images are resized to $3\times 384\times 384$. The CDIS is trained for 1 epoch, while the CDTS runs for 30 epochs, with a batch size of 64. In DBSCAN, the neighborhood radius is set to 0.40 and 0.30 for drone-view images and satellite-view images, respectively, and the minimum number required to form a cluster is 4. AdamW optimizer is used for CDIS with an initial learning rate of 0.001, followed by SGD optimization for the CDTS with a learning rate of 0.00025. Furthermore, in Eq.(7), $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 0.6, 0.1, and 1, respectively. The CDTS is activated only when there exist unpaired data.
IV-C Comparison with State-of-the-art Methods
We conduct experiments with three settings, namely (i) fully supervised learning, (ii) few-shot initialization, and (iii) cross-domain initialization. The methods in comparison encompass a range of state-of-the-art techniques. For supervised learning, these include MuSe-Net [45], TransFG [56], IFSs [15], MCCG [36], Sample4Geo [9], CAMP [47] and DAC [48]. In the realm of unsupervised learning, the methods are EM-CVGL [24], SegVLAD [14], AnyLoc [21] and Li [23]. Notably, SegVLAD [14], AnyLoc [21] and Li [23] are reported in separate Tables IV, V and VI due to the absence of complete results in their publications.
[…]
TABLE I: Comparisons between existing state-of-the-art methods and CDIKTNet on University-1652.1111
University-1652
[…]
R@1 AP R@1 AP
i MuSe-Net[45] 100 74.48 77.83 88.02 75.10
TransFG[56] 100 84.01 86.31 90.16 84.61
IFSs[15] 100 86.06 88.08 91.44 85.73
MCCG[36] 100 89.40 91.07 95.01 89.93
Sample4Geo[9] 100 92.65 93.81 95.14 91.39
CAMP[47] 100 94.46 95.38 96.15 92.72
DAC[48] 100 94.67 95.50 96.43 93.79
CDIKTNet 100 95.02 95.79 96.15 93.93
$i$ CDIKTNet 10 88.02 89.84 93.87 86.48
CDIKTNet 2 85.63 87.75 90.44 82.24
UM EM-CVGL[24] 0 70.29 74.93 79.03 61.03
iii CDIKTNet 0 83.30 85.73 87.73 76.53
“UM” is employed as the abbreviated form of “Unsupervised Method”.
TABLE II: Comparisons between existing state-of-the-art methods and CDIKTNet on SUES-200.
SUES-200
[…]
150m 200m 250m 300m 150m 200m 250m 300m
R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP
i IFSs [15] 100 77.57 81.30 89.50 91.40 92.58 94.21 97.40 97.92 93.75 79.49 97.50 90.52 97.50 96.03 100.00 97.66
MCCG[36] 100 82.22 85.47 89.38 91.41 93.82 95.04 95.07 96.20 93.75 89.72 93.75 92.21 96.25 96.14 98.75 96.64
Sample4Geo[9] 100 92.60 94.00 97.38 97.81 98.28 98.64 99.18 99.36 97.50 93.63 98.75 96.70 98.75 98.28 98.75 98.05
DAC[48] 100 96.80 97.54 97.48 97.97 98.20 98.62 97.58 98.14 97.50 94.06 98.75 96.66 98.75 98.09 98.75 97.87
CAMP[47] 100 95.40 96.38 97.63 98.16 98.05 98.45 99.33 99.46 96.25 93.69 97.50 96.76 98.75 98.10 100.00 98.85
CDIKTNet 100 96.03 96.87 97.80 98.27 98.90 99.15 99.73 99.79 97.50 94.23 98.75 96.58 100.0 98.33 98.75 98.84
$i$ CDIKTNet 10 83.60 86.24 90.53 92.14 94.33 95.32 97.13 97.60 91.25 83.24 95.00 89.53 97.50 93.59 97.50 95.80
CDIKTNet 2 80.55 83.58 88.05 90.10 91.28 93.13 94.48 96.12 90.00 77.16 91.25 85.92 97.50 91.36 97.50 94.07
UM EM-CVGL[24] 0 55.23 60.80 60.95 61.03 68.10 72.62 74.42 78.20 73.75 54.00 91.25 65.65 96.25 72.02 97.50 74.74
iii CDIKTNet 0 82.75 85.25 89.35 91.13 93.15 94.39 95.18 96.12 88.75 80.33 93.75 88.17 95.00 92.15 98.75 94.37
TABLE III: Comparisons between existing state-of-the-art methods and CDIKTNet trained in setting i and $i$ in generalization capacity.
[…]
150m 200m 250m 300m 150m 200m 250m 300m
R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP
i MCCG[36] 100 57.62 62.80 66.83 71.60 74.25 78.35 82.55 85.27 61.25 53.51 82.50 67.06 81.25 74.99 87.50 80.20
Sample4Geo[9] 100 70.05 74.93 80.68 83.90 87.35 89.72 90.03 91.91 83.75 73.83 91.25 83.42 93.75 89.07 93.75 90.66
DAC[48] 100 76.65 80.56 86.45 89.00 92.95 94.18 94.53 95.45 87.50 79.87 96.25 88.98 95.00 92.81 96.25 94.00
CAMP[47] 100 78.90 82.38 86.83 89.28 91.95 93.63 95.68 96.65 87.50 78.98 95.00 87.05 95.00 91.05 96.25 93.44
CDIKTNet 100 80.20 81.20 88.85 90.65 93.32 94.51 94.10 95.16 90.00 82.85 93.75 89.92 96.25 92.48 95.00 93.62
$i$ CDIKTNet 10 81.35 84.28 89.15 91.16 94.18 95.29 96.68 97.32 91.25 81.10 97.50 90.44 98.75 93.87 98.75 95.05
CDIKTNet 2 70.98 75.15 80.18 83.47 87.05 89.36 87.05 94.37 81.25 71.05 92.50 81.20 92.50 87.60 96.25 91.49
UM EM-CVGL[24] 0 60.03 65.69 71.50 76.17 80.35 83.85 85.93 88.34 73.75 56.99 87.50 70.62 92.50 81.18 95.00 86.04
iii CDIKTNet 0 79.08 82.55 90.07 91.85 93.05 94.37 95.00 95.95 86.25 77.90 92.50 86.40 93.75 90.16 95.00 92.22
TABLE IV: Comparisons between existing state-of-the-art unsupervised methods and CDIKTNet on university-1652.2222
University-1652
Setting Model GT Ratio (%) Drone→Satellite
R@1 R@5 R@10 AP
UM AnyLoc[21] 0 42.08 57.92 69.19 -
SegVLAD[14] 0 68.19 77.32 77.18 -
Li(a)[23] 0 57.49 73.32 77.18 -
Li(b)[23] 0 76.60 80.46 84.88 -
$i$ CDIKTNet 10 88.02 89.84 93.87 86.48
CDIKTNet 2 85.63 94.86 96.31 87.75
iii CDIKTNet 0 83.30 94.00 95.71 85.73
Li(a)[23] and Li(b)[23] correspond to different parameter configurations reported in the original paper.
TABLE V: Comparisons between existing state-of-the-art unsupervised methods and CDIKTNet on SUES-200
SUES-200
Setting Model GT Ratio (%) Drone→Satellite
200m 300m
R@1 AP R@1 AP
UM AnyLoc[21] 0 34.00 46.74 41.00 51.94
SegVLAD[14] 0 69.50 72.98 73.00 74.26
Li(a)[23] 0 56.00 60.50 62.00 65.00
Li(b)[23] 0 73.00 76.54 76.50 77.52
$i$ CDIKTNet 10 89.15 91.16 96.68 97.32
CDIKTNet 2 88.05 90.10 94.48 96.12
iii CDIKTNet 0 89.35 91.13 95.18 96.12
TABLE VI: Comparisons between existing state-of-the-art unsupervised methods and CDIKTNet trained in setting ii in generalization capacity.
[…]
Setting Model GT Ratio (%) Drone→Satellite
150m 200m 250m 300m
R@1 AP R@1 AP R@1 AP R@1 AP
UM AnyLoc[21] 0 30.00 42.74 34.00 46.74 36.78 50.41 41.00 51.97
SegVLAD[14] 0 56.00 59.08 69.50 72.98 72.50 74.10 73.00 74.26
Li(b)[23] 0 66.00 69.98 73.00 76.54 73.50 75.15 76.50 77.52
$i$ CDIKTNet 10 81.35 84.28 89.15 91.16 94.18 95.29 96.68 97.32
CDIKTNet 2 70.98 75.15 80.18 83.47 87.05 89.36 87.05 94.37
iii CDIKTNet 0 79.08 82.55 90.07 91.85 93.05 94.37 95.00 95.95
TABLE VII: Comparisons between existing state-of-the-art methods and CDIKTNet on SUES-200 with 2% and 10% paired data.3333
SUES-200
[…]
150m 200m 250m 300m 150m 200m 300m 300m
R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP
- Sample4Geo[9] 2 62.10 68.11 69.85 74.93 74.85 79.14 76.40 80.33 77.50 57.91 88.75 70.17 91.25 76.99 90.00 78.89
CAMP[47] 2 54.73 61.70 62.30 68.39 65.83 71.42 68.40 73.58 68.75 49.25 82.50 60.49 87.50 68.79 91.25 74.43
DAC[48] 2 55.83 62.16 68.50 73.25 73.38 77.34 74.70 78.64 70.00 51.09 85.00 61.59 90.00 70.53 91.25 74.63
$i$ CDIKTNet 2 80.55 83.58 88.05 90.10 91.28 93.13 94.48 96.12 90.00 77.16 91.25 85.92 97.50 91.36 97.50 94.07
- Sample4Geo[9] 10 75.43 79.21 84.20 86.70 87.18 89.40 89.63 91.61 91.25 76.89 92.50 82.76 96.25 86.60 96.25 89.23
CAMP[47] 10 76.27 80.39 83.83 86.89 89.66 91.62 90.85 92.71 92.50 77.50 96.25 86.63 97.50 91.45 97.50 93.29
DAC[48] 10 76.80 80.65 85.05 87.65 88.80 90.81 90.58 92.32 93.75 79.06 97.50 86.78 97.50 90.58 98.75 92.54
$i$ CDIKTNet 10 83.60 86.24 90.53 92.14 94.33 95.32 97.13 97.60 91.25 83.24 95.00 89.52 97.50 93.59 97.50 95.80
“–” denotes the scenario where supervised methods are applied under limited paired relationships.
TABLE VIII: Comparisons between existing state-of-the-art methods and CDIKTNet on University-1652 with 2% and 10% of paired data.
University-1652
[…]
R@1 AP R@1 AP
- Sample4Geo[9] 2 62.41 66.85 84.45 60.52
CAMP[47] 2 62.53 67,01 83.31 59.26
DAC[48] 2 63.68 68.16 87.59 61.26
$i$ CDIKTNet 2 85.63 87.75 90.44 82.24
- CAMP[47] 10 83.64 86.27 92.44 82.27
Sample4Geo[9] 10 84.51 86.94 93.01 83.53
DAC[48] 10 84.64 87.07 93.15 83.03
$i$ CDIKTNet 10 88.02 89.84 93.87 86.48
(ii) Few-shot Initialization: To evaluate the effectiveness of CDIKTNet in leveraging unlabeled data for learning, we initialize CDIKTNet with only 2% and 10% of the paired data. The remaining training data are entirely unpaired. As shown in Tables I and II, CDIKTNet with only 2% of the paired data achieves performance comparable to or even surpassing that of IFSs [56]. With 10% of paired data, CDIKTNet achieves further improvements, surpassing most fully supervised methods and achieving performance comparable to MCCG [36]. Notably, as shown in Table III, CDIKTNet in this setting trained with only 10% of the paired data exhibits even stronger generalization than these fully supervised methods.
Furthermore, compared with state-of-the-art unsupervised methods, including EM-CVGL [24], AnyLoc [21], SegVLAD [14], and Li [23], as illustrated in Tables I, II, IV and V, CDIKTNet in this setting significantly outperforms them in all metrics on both datasets, despite relying on only a small proportion of paired data. In terms of generalization, CDIKTNet still significantly surpasses these methods, as reported in Tables III and VI. This underscores the significance of employing a limited quantity (e.g., 2%) of paired data for initialization, as it serves to steer the exploration toward revealing deeper cross-view representations from unpaired data.
(iii) Cross-domain Initialization: A novel experimental setting is proposed for cross-domain validation. CDIKTNet is initialized with 2% of paired data on one dateset and then conducts the training process with the remaining unpaired data, whose whole samples that corresponds to a specific domain. This trained model is then transferred to another dataset representing a distinct domain for further training without requiring additional pairing relationships. Notably, this transfer learning phase operates with 0% of paired data. Therefore, in this setting, the GT ratio is marked as 0%.
In setting iii of Tables I, and IV, CDIKTNet is first initialized by using 2% paired data on SUES-200 dataset and then continue training with the remaining unpaired data. The resulting model has corresponded to setting ii with GT Ratio 2% in Table II. This model is subsequently used as a cross-domain initialization and transferred to the other dataset University-1652 for further training, where no paired data are required in the following process. Similarly, in setting iii of Tables II and V, CDIKTNet is initialized using 2% paired data on the University-1652 dataset and continue training with its remaining unpaired data. This trained model has reported in setting ii with GT Ratio 2% in Table I. This model is then used as a cross-domain initialization on the SUES-200 dataset, where the transfer stage also relies entirely on unpaired data on this new dataset. Furthermore, the results in setting iii of Tables III and VI are obtained by directly evaluating the models from setting iii of Tables I and IV on the SUES-200 dataset, without requiring any additional fine-tuning.
As shown in Tables I, II, IV and V, after training in one domain, CDIKTNet can be directly transferred to another domain. It outperforms state-of-the-art unsupervised methods in all metrics and achieves performance comparable to several fully supervised methods. Moreover, as reported in Tables III and VI, CDIKTNet demonstrates remarkable generalization performance. It significantly outperforms these state-of-the-art unsupervised methods. Notably, as shown in V, the experimental results in this setting demonstrates superior performance compared with the few-shot initialization setting, which has utilized 2% of paired data in current domain. This indicates that strong domain adaptability enables the cross-domain initialization setting to outperform limited paired supervision under some cases.
Remark 4
The (ii) few-shot initialization setting and (iii) cross-domain initialization setting are two typical scenarios of limited supervision. The former targets the current domain by using a limited amount of paired data for initialization, which is a process of supervised learning. Subsequently, unpaired data is employed for unsupervised learning. The latter addresses the issue of a complete lack of paired data in a new domain. First, it performs initialization through supervised learning in an existing domain using a limited amount of paired data. Then, it uses the remaining unpaired data for unsupervised learning. The obtained model is directly applied to unsupervised learning in the new domain, and has achieved remarkable results in the DVGL task in the new domain.
Furthermore, as shown in Tables VII and VIII, existing supervised methods strictly rely on paired data as positive samples and cannot use unpaired data. In contrast, CDIKTNet can effectively exploit a large amount of unpaired cross-view information to achieve better performance under such limited conditions. Additionally, CDIKTNet also achieves competitive performance with 100% of paired data equivalent to fully supervised learning. This highlights its excellent adaptability and generalization in different supervision levels.
TABLE IX: Influence of each component on performance of CDIKTNet.
University-1652
[…]
R@1 AP R@1 AP
2% 58.19 62.69 82.45 56.95
$\checkmark$ 59.00 63.55 81.16 55.03
$\checkmark$ 59.73 64.16 80.03 57.35
$\checkmark$ $\checkmark$ 63.93 68.19 83.45 61.35
$\checkmark$ $\checkmark$ $\checkmark$ 85.63 87.75 90.44 82.24
100% 92.80 94.00 95.01 91.83
$\checkmark$ 94.02 94.96 95.72 93.16
$\checkmark$ 93.26 94.36 95.58 92.15
$\checkmark$ $\checkmark$ 95.02 95.79 96.15 93.93
IV-D Ablation Studies
Effectiveness of CDIS. As shown in Table IX, both the SIFL block and SIEL module contribute significantly to performance improvements with both 2% and 100% of paired data. When used individually, each component increases R@1 and AP, which demonstrates their individual effectiveness. Notably, their joint optimization yields further improvement, which validates the design of CDIS. Specifically, SIFL enhances geometric invariance at different viewpoints by modeling structural priors, while SIEL improves spatial invariance by incorporating scale-aware contextual cues. These two components complement each other to enable CDIKTNet to construct a unified and robust cross-view shared feature space.
[…]
IV-E Further Analysis
Hyper-parameters Analysis of Loss Coefficients. We further investigate the influence of the loss coefficients $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ in Eq.  (7) on the University-1652 dataset. As shown in Fig. 4, the performance influence of all three hyper-parameters exhibit bounded fluctuations within reasonable ranges. It achieves consistently robust performance for $\lambda_{1}\in[0.4,0.8]$, and values outside this range diminish the effectiveness. Then the peak performance occurs near $\lambda_{2}=0.10$, with marginal degradation observed at both lower and higher values. Furthermore, it exhibits a gradual improvement as its weight increases from 0.40 to 1.00 for $\lambda_{3}$, in which the best results are obtained when $\lambda_{3}=1.00$. Synthesizing these observations, the highest overall performance is attained at $\lambda_{1}=0.60$, $\lambda_{2}=0.10$, and $\lambda_{3}=1.00$. Consequently, these coefficient values are adopted for all subsequent experiments.
Figure 4: Performance Influence of $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$. Influence of different $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ values on University-1652 in setting iii.
[…]
Figure 5: Performance Influence of $\alpha$. Influence of different $\alpha$ values on University-1652 in setting iii.
Figure 6: Influence of CDIS on Convergence, Performance, and Training Efficiency. (a) Loss curves and statistical summaries for baseline vs. baseline with CDIS on University-1652 in setting i. (b) Comparison of performance and training time between baseline and baseline with CDIS on University-1652 in setting i.
Influence of CDIS on Convergence, Performance, and Training Efficiency. To further evaluate the influence of CDIS on convergence, the training behaviors of the baseline and baseline with CDIS are compared on the University-1652 dataset in setting i. As shown in Fig. 6(a), it presents the loss curves and their statistical values. The baseline and baseline with CDIS exhibit similar descending trends and stable convergence regions. The baseline with CDIS begins with a slightly higher loss but reaches a lower final value, which demonstrates that the optimization process remains stable and that CDIS guides the model toward a better solution. Fig. 6(b) reports the performance and the training time. The CDIS improves the performance in both the drone $\rightarrow$ satellite and satellite $\rightarrow$ drone scenarios compared to the baseline, and the additional training cost is small at about 35 seconds. Empirical evidence confirms that CDIS simultaneously enhances structural and spatial invariance, which maintains rapid convergence rates and incurs only marginal computational overhead, ultimately achieving superior performance.
Figure 7: Similarity Distribution and Feature Space Visualization on University-1652 . (a–e) show the similarity distribution visualization. The red-shaded area indicates the overlapping region between the positive and negative distributions. (f-j) show the distribution of feature embeddings in the 2D shared feature space, where circles and triangles in different colors denote drone-view and satellite-view. A total of 20 localizations are randomly selected from the test set.
Figure 8: Attention Heatmap Visualization. The attention responses generated by the proposed SIEL module are visualized. Red regions correspond to stronger attention responses, whereas blue regions denote weaker responses.
Figure 9: Retrieval Visualization on University-1652. The green box indicates a correctly-matched image, and the red box indicates a falsely-matched image.
TABLE X: Comparisons between existing state-of-the-art methods and CDIKTNet on DenseUAV
DenseUAV
All height 80m 90m 100m
Setting Model GT Ratio (%) R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5
i ConvNext-Tiny[7] 100 60.23 81.94 - - - - - -
Sample4Geo[9] 100 80.57 96.53 80.57 96.53 86.87 98.71 91.38 99.74
DAC[48] 100 84.47 96.53 80.92 97.04 85.32 98.71 85.46 99.10
CDIKTNet 100 85.50 97.53 83.14 97.55 86.36 99.23 87.39 99.49
$i$ CDIKTNet 10 51.57 79.37 49.93 83.91 52.90 84.93 52.12 45.97
CDIKTNet 2 39.76 72.11 39.00 75.55 39.64 78.77 39.90 80.31
UM EM-CVGL 0 18.15 50.97 18.02 50.36 17.76 50.19 17.50 49.98
iii CDIKTNet 0 22.57 40.24 17.25 40.15 23.04 48.39 25.48 52.38
[…]
IV-F Visualization
As shown in Fig. 7, to further demonstrate the effectiveness of CDIKTNet, we visualize the similarity distribution using t-SNE [44] and feature space projection.
Similarity Distribution Visualization. As shown in Fig. 7(a–b), Sample4Geo and DAC achieve overlap areas of 0.1232 and 0.1124 under limited supervision, due to their inability to exploit unpaired data. Fig. 7(c) shows that EM-CVGL [24] achieves an overlap of 0.1158. Although it performs slightly better, feature confusion remains an issue. In contrast, CDIKTNet in Fig. 7(d) achieves a lower overlap of 0.0661 with only 2% of paired data. This confirms its strong ability to align and separate cross-view features under limited supervision. Notably, in Fig. 7(e), CDIKTNet is trained in the cross-domain initialization setting described in Subsection IV-C. It still achieves an overlap of 0.0808 to outperform all methods. This shows its effectiveness in exploiting unpaired data and transferring knowledge from an external source domain.
Feature Space Visualization. In the feature space, features from the same location should cluster, while different locations should be well separated. As shown in Fig. 7(f–h), the compared methods show noticeable cross-view confusion (red circles) and dispersed feature distributions. In contrast, CDIKTNet (Fig. 7(i–j)) produces more compact and structured feature distributions in both scenarios. Although the 0% of paired data in cross-domain initialization setting shows minor confusion, the features remain more compact and more evenly aligned. This presents the ability of CDIKTNet to learn cross-view information.
Attention Heatmap Visualization. As a further demonstration of the effectiveness of SIEL, we visualize its attention heat maps in Fig. 8. Despite the large spatial differences between satellite-view and drone-view images and the feature distribution shifts caused by varying drone viewpoints, SIEL consistently highlights semantically meaningful regions in both views.
[…]
V Conclusion
This paper addresses a critical and highly challenging problem in the drone-view geo-localization (DVGL) task with minimal or even no paired data to reduce dependence on expensive cross-view annotations. To this end, we propose a novel cross-domain invariant knowledge transfer network (CDIKTNet), which learns the spatial and structural invariance of cross-view features while constructing a robust invariant knowledge representation space. Meanwhile, it utilizes this shared space as an optimization anchor for unsupervised training, which effectively learns latent cross-view information from unpaired data. It significantly alleviates the reliance on paired data. More importantly, CDIKTNet can be applied under any level of supervision and can be generalized to new environments without requiring their prior pairing relationships. Extensive experiments demonstrate that, under full supervision, CDIKTNet outperforms existing state-of-the-art methods. Furthermore, when initialized with a small amount of paired data in the current domain or transferred to the other domain in the proposed cross-domain initialization setting, CDIKTNet outperforms state-of-the-art unsupervised methods and even surpasses some fully supervised methods. This advancement pushes DVGL closer to real-world deployment.